OpenStudy (anonymous):

Find domain of : F(x)= square root of 2x + 6 over 4x^2 -4x -3

OpenStudy (anonymous):

You know that the \[\sqrt{2x+6}\] must be greater than 0 therefore x>=3 you must realize the bottom also must be greater than 0 since we cannot divide by 0. try setting 4x^2-4x-3>=0 =>4x^2-4x>=3 factor out a 4x and solve. Hope this helps. Gives you some suggestion.